7 3/5
- 4 5/6
_______________
?

Answer by checkley77(12844)

7 3/5 OR 38/5 MULT BY 6=228/30
-4 5/6 OR -29/6 MULT BY 5=-145/30
228/30-145/30=83/30=2 23/30=2.76666667